OpenStudy (anonymous):

Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). The two points are (63,63) and (72,70)

OpenStudy (anonymous):

I need help understanding this, cause I've tried different times to get an answer finding the values but it never comes out right... this is what I got: y-63=7/9(x-63) y-63=7/9x+39.69 +63 +63 Y=7/9x+102.69

OpenStudy (anonymous):

I had to make a scatter plot graph with y= people's height, and x= arm span (all in inches)

OpenStudy (harsimran_hs4):

7/9 * 63 = 49 perhaps not 39.69

OpenStudy (anonymous):

So is the equation I wrote out alright (besides the 49)? and yeah I started working it out different ways on a piece of paper and I got 49 as well, though the results come out strange, because when I multiply 49 by the input x value, it gives me an outrageous number :| like 3k something, nobody's got arms THAT long...

OpenStudy (harsimran_hs4):

also it should be -49(you missed the negative sign) there instead of 39.49

OpenStudy (anonymous):

ohhh right, it's negative, so it's y-63=7/9x-49 ? but don't I have to multiply 49*63? that gives me 3087.... cause I'm supposed to distribute the 7/9 which .. if I got rid of the fraction it would be 63.... so.... that's the part I'm stuck on

OpenStudy (harsimran_hs4):

no ... Y - 63=7/9x - 49 y = 7/9x -49 +63 y = 7/9x +14

OpenStudy (anonymous):

There we go!! Yeah I just solved it now as well, I realized I was distrributing incorrectly. Thank you! So one last question, if I wanted to input a new y-value, can I use this exact equation to find the x value?

OpenStudy (harsimran_hs4):

yes you can do that you can put any one of the x or y and obtain y or x respectively

OpenStudy (anonymous):

I tried to solve it and find the value of x using the slope-intercept equation--> y=7/9x+14 , so I substituted the y for 66, subtracted 14 from both sides, got 52=7/9x and that's where I get stuck, do I multiply the denominator to both sides ? cause I'd need to isolate x, right?

OpenStudy (harsimran_hs4):

multiply both sides by 9(denominator)

OpenStudy (anonymous):

Okay I got 468=63x , so it got canceled out on the right side, but on the left side do I solve it? it gives b=me 7.4 ... which must be wrong... because no one can be 7.4 inches short

OpenStudy (harsimran_hs4):

(7/9) * 9 = 7 not 63

OpenStudy (anonymous):

oohh.. I learned it as (9)7/9 the denominator cancels out and it's 7*9.... but right, it explains why I'm getting the wrong answer... I must've not realized it while learning about it. Thank you very much for all your help by the way!! :)

OpenStudy (harsimran_hs4):

Welcome :)